OpenStudy (anonymous):

i need help solving : x^2-12x+36=24

OpenStudy (anonymous):

First, get everything on the left side of the equation so it all is = 0 on the right. In other words, subtract 24 from both sides...

OpenStudy (anonymous):

okay so x^2-12x+12=0 do i now factor it?

OpenStudy (anonymous):

well, maybe that will help... I was thinking you could factor easily, but I was wrong. Do you know the quadratic formula or "completing the square"?

OpenStudy (anonymous):

quad=\[[-b +/- \sqrt{-b ^{2-4ac}}\] alll over 2a

OpenStudy (anonymous):

yes, so you can use that with the equation: x^2-12x+12 = 0 a = 1, b = -12, and c = 12 You get two results, one for each of the plus and minus on the quad formula. Those are the points where this this expression is true... the solutions for x.

OpenStudy (anonymous):

i got \[\frac{ 12+/-\sqrt{\sqrt{0}}}{ 2 }\]

OpenStudy (anonymous):

-b = 12, so the first term is right, and 2a = 2*1 = 2, so the denominator is right. inside the sqrt, it should be b^2 - 4ac = (-12)^2 - 4(1)(12) = 144 - 48 = 96 sqrt(96) simplifies to sqrt(16*6) = 4sqrt(6)

OpenStudy (anonymous):

oh i had put the original in sorry. then can i reduce by the 2 ??

OpenStudy (anonymous):

so I get\[\frac{ 12 \pm 4\sqrt{6}}{ 2 }\]

OpenStudy (anonymous):

i got 6+or-2root6

OpenStudy (anonymous):

sqrt(144 - 48) = sqrt(96) = sqrt(16 * 6) = 4sqrt(6)

OpenStudy (anonymous):

so it would be 6 + or - 3sqrt(6) after you cancel the 2 on the bottom with the 12 and 6 on top.

OpenStudy (anonymous):

sorry, typo above... should have been 2 before the sqrt, not 3 6 + or - 2sqrt(6) after you cancel the 2 on the bottom with the 12 and 6 on top.

OpenStudy (anonymous):

that's what you got, I guess :) Good work... I kept making small math and typing errors...

OpenStudy (anonymous):

okay thanks. when do i know to do quadratic formula and completing the square. because i forgot how to do completing the square.

OpenStudy (anonymous):

I never remember completing the square! You can always use quadratic formula if you are good at it. However, if factoring works, then that is easier. But if you can't factor, and you can't remember completing the square, feel free to use quadratic formula. The only exception is if you will be tested specifically on completing the square... if so, learn it for the test :) It's a good method, but it isn't the only way.

OpenStudy (anonymous):

okay thanks !!!

OpenStudy (anonymous):

glad to help... sorry I confused myself for a minute there... :)